Lisa​'s Floral​ Arrangements, Inc. had the following transactions in the month of​ January: The owners invested $90,000 ​(the par value of the​ stock) for​ 20,000 shares of common​ stock; the company purchased furniture for the florist shop in the amount of $35,000​, which was put on a credit account with the​ vendor; and the company paid employees wages of $14,000 in cash. What is the effect of each of these transactions on the accounting​ equation?

Cash + Furniture = Liabilities + Contributed Capital + Beg. Retained Earnings + Revenues & Gains - Expenses & Losses - Dividends Declared
Answer #1
Cash + Furniture = Liabilities + Contributed Capital + Beg. Retained Earnings + Revenues & Gains - Expenses & Losses - Dividend Declared
1 90,000 + = + 90,000 + + - -
2 + 35,000 = 35,000 + + + - -
3 -14,000 + = + + + - -14,000 -
Total 76,000 + 35,000 = 35,000 + 90,000 + 0 + 0 - -14,000 - 0

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks